# From Equations to Insights: Unraveling Symbolic Structures in PDEs with LLMs

### v1

@@ Abstract
Motivated by the remarkable success of artificial intelligence (AI) across diverse fields, the application of AI to solve scientific problems—often formulated as partial differential equations (PDEs)—has garnered increasing attention. While most existing research concentrates on theoretical properties (such as well-posedness, regularity, and continuity) of the solutions, alongside direct AI-driven methods for solving PDEs, the challenge of uncovering symbolic relationships within these equations remains largely unexplored. In this paper, we propose leveraging large language models (LLMs) to learn such symbolic relationships. Our results demonstrate that LLMs can effectively predict the operators involved in PDE solutions by utilizing the symbolic information in the PDEs. Furthermore, we show that discovering these symbolic relationships can substantially improve both the efficiency and accuracy of the finite expression method for finding analytical approximation of PDE solutions, delivering a fully interpretable solution pipeline. This work opens new avenues for understanding the symbolic structure of scientific problems and advancing their solution processes.
Keywords Large Language Models  $\cdot$ Finite Expression Method  $\cdot$ High Dimensional PDEs  $\cdot$ Operator Relation
@@ 1 Introduction
Numerous grand challenges in scientific fields, including physics, engineering, biology, and chemistry, can be modeled as partial differential equations (PDEs), which relate unknown functions defined on a multidimensional domain (e.g., in terms of space and time) to its rates of change with respect to each of those variables (Evans, 2022). Mathematically, a PDE takes the following form
$Du(x)=f(u(x),x),\quad x\in\Omega,\quad Bu(x)=g(x),\quad x\in\partial\Omega,$ (1)
where $D$ denotes a certain differential operator defining the dynamics, $Bu=g$ represents the boundary condition (such as Dirichlet, Neumann, and Robin), and $\Omega\subseteq\mathbb{R}^{d}$ describes the domain with boundary $\partial\Omega$.
PDEs often arise in high-dimensional settings (i.e., $d$ is large); for example, well-known instances such as the Poisson equation (Yu et al., 2018), the linear conservation law (Chen et al., 2021), and the nonlinear Schrödinger equation (Han et al., 2020) regularly involve multiple spatial and temporal variables. While low-dimensional PDEs can be solved reliably and efficiently via traditional mesh-dependent methods such as finite difference methods (LeVeque, 2007), finite elements methods (Hughes, 2003), finite volume methods (Toro, 2013), and spectral methods (Boyd, 2001), developing efficient and accurate algorithmic frameworks for computing numerical solutions to high-dimensional PDEs remains an important and challenging topic due to the curse of dimensionality (i.e., the computational costs grow exponentially with respect to the dimensionality of the problem’s domain) (E et al., 2021).
Before solving a PDE, one must understand certain fundamental properties of the solution $u(x)$ that follow from the PDE itself and from the data—namely, $f(u(x),x)$ and $g(x)$. These considerations align with the classical theory of PDEs (Evans, 2022), which focuses on establishing well-posedness (i.e., the existence and uniqueness of solutions) and their regularity and continuity. However, the underlying symbolic relationship among $u$, $f$ and $g$ often remains underexplored. To address this gap, we pose the following question:
Assuming that the PDE (1) admits an analytical solution $u$, and given the operators appearing in the expressions of $f$ and $g$, what are the corresponding operators in the expression for $u$?
Understanding this symbolic connection is crucial for deriving fully interpretable, closed-form solutions, a feature lacking in most existing solution methods, particularly in the high-dimensional setting. As an illustrative example, we consider solving the PDE (1) via the least square method (Dissanayake and Phan-Thien, 1994; Lagaris et al., 1998; Sirignano and Spiliopoulos, 2018), which defines a straightforward functional to characterize the error of the estimated solution by
@@
$u^{*}=\mathrm{argmin}_{u\in L^{2}(\Omega)}\;L(u).$ (3)
Because the search space $L^{2}(\Omega)$ is an infinite dimensional space, finding the solution $u^{*}$ is extremely challenging. Fortunately, the search space can be significantly simplified if one knows the operators appearing in the solution in advance, making the search of the optimal solution much easier and effective. This can be accomplished by adopting the recently developed finite expression method (FEX) (Liang and Yang, 2022).
To uncover the symbolic relationship described above, we leverage large language models (LLMs) as predictive tools. The process begins with the generation of a comprehensive, structured dataset composed of symbolic expressions derived from a diverse range of PDE types, including elliptic, parabolic, and hyperbolic equations, along with their associated boundary and initial conditions. These symbolic expressions are represented as computational trees (see Figure 1), where nodes are populated with unary (e.g., sin, cos, exp) and binary operators (e.g., +, -, *, /). This tree structure captures the hierarchical nature of mathematical expressions and ensures a systematic representation of the PDE solutions.
@@ 1.1 Related work
Artificial intelligence (AI) for Science takes advantage of modern advanced machine learning (ML) methods to facilitate, accelerate and enhance scientific discovery (Stevens et al., 2020). Its importance lies in the ability to handle massive datasets, uncover hidden patterns, and simulate complex phenomena. Inspired by the remarkable success of AI across various domains, deep neural network (DNN)-based approaches (LeCun et al., 2015; Goodfellow, 2016) have become increasingly popular for solving (high-dimensional) PDEs. These methods offer certain appealing advantages over traditional numerical techniques, making them a compelling alternative for tackling complex problems. First, DNN-based approaches empirically mitigate the curse of dimensionality that plagues conventional discretization approaches, leading to more efficiency in handling high-dimensional PDEs Han et al. (2018). Second, DNNs, acting as mesh-free universal function approximators (Cybenko, 1989; Hornik et al., 1989; Goodfellow, 2016), can be trained to learn highly complex PDE solutions without the need of prespecifying basis functions or meshes (Berg and Nyström, 2018; Raissi et al., 2019). Hence, they are highly suitable to approximate complex solution spaces that challenges the traditional mesh-based approaches. Third, the solutions represented by DNNs can be evaluated efficiently once trained, which turns out to be a crucial property for tasks, including uncertainty quantification and real-time optimal control, require calling PDE solutions repeatedly (Zhu et al., 2019). Last but not least, continuous progress in designing comprehensive neural network architectures—such as Fourier neural operators and advanced physics-informed neural network frameworks—alongside theoretical advancements in understanding their approximation capabilities, are making these methods increasingly applicable (Lu et al., 2019; Li et al., 2020).
@@
However, several critical factors often hinder deep neural networks (DNNs) from achieving highly accurate solutions, even for relatively simple problems. These include the need for large, diverse, and high-quality training data (Szegedy, 2013; Buda et al., 2018), sensitivity to hyperparameter selection (Bergstra and Bengio, 2012), the challenge of optimizing highly nonconvex objectives (Dauphin et al., 2014; Choromanska et al., 2015), issues like vanishing and exploding gradients (Glorot and Bengio, 2010), and poor generalization capabilities (Zhang et al., 2021). Additionally, solutions produced by DNNs often lack interpretability, preventing users from leveraging insights for future decision-making (Rudin, 2019). To overcome these challenges and achieve both highly accurate and interpretable solutions for PDEs, Liang and Yang (2022) recently introduced the finite expression method (FEX). This approach leverages recent advances in approximation theory, which demonstrate that high-dimensional functions can be effectively approximated using functions composed of simple operators (Shen et al., 2021a, b; Zhang et al., 2022). This methodology seeks approximate PDE solutions within the space of functions composed of finitely many analytic expressions, effectively avoiding the curse of dimensionality. In FEX, the mathematical expression representing the PDE solution is modeled as a binary tree, where each node holds either a unary or binary operator chosen from pre-specfied operator sets. The goal is to find the optimal sequence of operators and coupling parameters that minimize the objective function associated with the PDE. This leads to a combinatorial optimization (CO) problem, which is then solved by a deep reinforcement learning (RL) method with the policy gradient optimizer. The FEX outperforms certain existing numerical PDE solvers in terms of accuracy, interpretability and memory efficiency, making it a promising approach for solving high-dimensional PDEs and other complex problems; see e.g., (Song et al., 2023, 2024; Hardwick et al., 2024).
@@ 2 Methodology
2.1 Binary Computational Trees for PDE Expressions
Expressions in PDEs are represented as binary computational trees, as depicted in Figure 1. These trees provide a structured and hierarchical way to encode mathematical expressions, enabling efficient exploration and manipulation of the solution space. The tree generation process is systematic and begins by specifying the depth of the binary tree, which determines the complexity and granularity of the resulting expressions.
Figure 1: Computational expression tree (Liang and Yang, 2022).
The construction of the tree involves the following steps:
  1. 1.
    Depth Specification: The depth of the tree is predefined, controlling the number of hierarchical levels in the tree. A deeper tree allows for more complex expressions but increases computational cost.
  2. 2.
    Recursive Construction: At each level of depth, the tree is built recursively. Starting from the root node, each node is expanded by adding child nodes according to the rules defined for operators. This recursive process ensures a systematic exploration of the solution space, covering a wide range of possible expressions.
Each node in the computational tree represents an operator, which can be either a unary operator or a binary operator:
  • Unary Operators: These operators act on a single operand and include functions such as trigonometric functions ($\sin(\cdot)$, $\cos(\cdot)$, $\tan(\cdot)$), exponential functions ($\exp(\cdot)$), logarithmic functions ($\lg(\cdot)$, $\ln(\cdot)$), and other nonlinear transformations ($\sqrt{\cdot}$, ||$\cdot$|| ⋅ |, $(\cdot)^{k}$, $k\in\mathbb{Z}$). For each node with a unary operator, two scalars ($\alpha$ and $\beta$) are coupled with the node. These scalars enable the composition of the underlying operation with an affine function, resulting in expressions of the form $\alpha\cdot u(\cdot)+\beta$, where $u(\cdot)$ is the unary operation. This coupling ensures that the tree can represent both nonlinear transformations and linear adjustments, enhancing its expressive power.
  • Binary Operators: These operators require two operands and include standard arithmetic operations including addition ($+$), subtraction (--), multiplication ($\times$) and division (///). Binary operators combine two sub-expressions (left and right child nodes) to form more complex expressions. For example, a multiplication node might combine the results of two subtrees representing different terms in an expression.
The use of binary computational trees offers several advantages: (1) The tree structure provides a clear and interpretable representation of mathematical expressions, making it easier to analyze and manipulate them; (2) By recursively constructing trees of varying depths, the method ensures a thorough exploration of the solution space, capturing both simple and complex expressions; and (3) The inclusion of both unary and binary operators, along with affine transformations, allows the tree to represent a wide range of mathematical operations and relationships. By systematically generating and evaluating computational trees, it becomes possible to identify candidate expressions that best describe the underlying dynamics of a system. Additionally, the framework can be combined with advanced techniques, including optimization and reinforcement learning, to refine the scalars and operators, further improving the accuracy of the discovered models.
@@ 2.2 Data Generation for Equation Types
In this subsection, we describe the data generation process for constructing a structured dataset of symbolic expressions derived from various PDE types. The dataset is designed to capture the symbolic relationships between the solution of a PDE and the associated problem data, including the boundary conditions and right-hand-side functions. The data generation process is outlined in Figure 2.
@@
The construction of the one data point involves the following steps.
PDE Type Specification. We first specify the type of PDE (e.g., Poisson, linear conservation law, and Schrödinger) and the type of boundary conditions (e.g., Dirichlet, Neumann and Robin) based on the information of the operators $D$ and $B$. This results in the first component of the data point, which includes the PDE type and boundary condition type of the data point.
Randomly generated function $u(x)$. Next, we generate a random function $u(x)$ by constructing a binary computational tree with a specified depth, as described in Section 2.1. The tree is populated with random unary and binary operators drawn from prescribed operator sets that are rich enough. Consequently, we obtain the expression tree for the solution function $u(x)$, which is used to derive the second component of the data point, i.e., expressions of the right-hand-side function $f$ and the function $g$ defined on the boundary of the domain.
Postfix Representation. After obtaining the expressions of the functions $u(x)$, $f(u(x),x)$, and $g(x)$, we convert them into postfix notation. This conversion simplifies the tokenization process and aligns the expressions with the input format required by LLMs. The postfix notation also eliminates the need for parentheses, making the expressions more interpretable and easier to parse. More importantly, it reflects the relative positional relationship of operators in the expression, which is crucial for the LLMs to learn the symbolic relationships between the operators. In the end, we obtain the second component of the data point, which includes the operators used in the right-hand-side function $f(u(x),x)$, and the function $g(x)$. These input component are then concatenated to generate the input. Meanwhile, the label of the data set is obtained similarly by converting the expression of the function $u(x)$ into postfix notation. See Figure 3 for the detailed convertion pipeline and a simple illustrated example.
Figure 3: Pipeline for converting data into postfix notation and an illustrated example.
2.3 Fine-tuning
Inspired by the success of LLMs, we propose fine-tuning the sequence-to-sequence Transformer architecture that leverages cross-attention to incorporate PDE context when generating solution operators. This architecture is selected because it effectively captures long-range dependencies and the complex hierarchical structure inherent in PDE metadata, thereby learning an effective mapping between symbolic operator sequences.
We first preprocess and tokenize the dataset generated as in the previous subsection, and then fine-tune the LLM using the proceed data using the common practices in the literature. The whole pipeline is outlined in Figure 4 and the key components in the pipeline are discussed in detailed as follows.
Figure 4: Pipeline for tokenization and embedding of operator sequences for LLM fine-tuning
Tokenization. We begin by loading the PDE data in jsonl format, shuffling it, and splitting it into training and validation sets. Each data point is then processed by a textual embedding layer that converts operator symbols (e.g., sin, cos, exp) into tokens. To ensure a robust and consistent representation, we use HuggingFace’s  [^1] https://github.com/huggingface AutoTokenizer, which selects an appropriate tokenization scheme based on the underlying pre-trained model. The tokenizer pads sequences to a fixed length and incorporates learnable positional embeddings. Additionally, an attention mask is generated to differentiate between valid and padding tokens, facilitating self-attention computations. Finally, padding tokens in the label sequences are replaced with $-100$, the default ignore_index in PyTorch’s  [^2] https://pytorch.org/ nn.CrossEntropyLoss, ensuring they are ignored during loss computation. This approach standardizes the transformation of textual operator representations into numerical tokens for model training.
Architecture and Finetuning: Next, we select a pretrained LLM with a sequence-to-sequence Transformer architecture. We fine-tune this model using the HuggingFace Trainer API, which provides a suite of tools for efficient optimization. During fine-tuning, we minimize the token-wise cross-entropy loss, which penalizes incorrect token predictions, including misclassification of the End-of-Sequence (EOS) token. Backpropagation is performed to update the model’s parameters according to the optimization algorithm. We emphasize the use of teacher forcing  [^3] Teacher forcing is a training strategy where, at each time step, the decoder is provided with the actual ground-truth token instead of its own previous prediction. during fine-tuning. By supplying the decoder with the full target sequence, this approach ensures that the token-wise cross-entropy loss is computed over aligned positions, preventing error accumulation and reinforcing the correct sequential dependencies more effectively and robustly.

### v2

@@ Abstract
Motivated by the remarkable success of artificial intelligence (AI) across diverse fields, the application of AI to solve scientific problems, often formulated as partial differential equations (PDEs), has garnered increasing attention. While most existing research concentrates on theoretical properties (such as well-posedness, regularity, and continuity) of the solutions, alongside direct AI-driven methods for solving PDEs, the challenge of uncovering symbolic relationships within these equations remains largely unexplored. In this paper, we propose leveraging large language models (LLMs) to learn such symbolic relationships. Our results demonstrate that LLMs can effectively predict the operators involved in PDE solutions by utilizing the symbolic information in the PDEs both theoretically and numerically. Furthermore, we show that discovering these symbolic relationships can substantially improve both the efficiency and accuracy of symbolic machine learning for finding analytical approximation of PDE solutions, delivering a fully interpretable solution pipeline. This work opens new avenues for understanding the symbolic structure of scientific problems and advancing their solution processes.
@@ 1 Introduction
Numerous grand challenges in scientific fields, including physics, engineering, biology, and chemistry, can be modeled as partial differential equations (PDEs), which relate unknown functions defined on a multidimensional domain (e.g., in terms of space and time) to its rates of change with respect to each of those variables [11]. Mathematically, a PDE takes the following form
$Du=f,\quad\text{in }\Omega,\quad Bu=g,\quad\text{on }\partial\Omega,$ (1)
where $D$ denotes a differential operator defining the dynamics, $Bu=g$ represents the boundary condition (e.g., Dirichlet, Neumann, and Robin), and $\Omega\subseteq\mathbb{R}^{d}$ is the domain with boundary $\partial\Omega$.
PDEs often arise in high-dimensional settings (i.e., $d$ is large); for example, well-known instances such as the Poisson equation [43], the linear conservation law [5], and the nonlinear Schrödinger equation [17] naturally involve multiple spatial and temporal variables. While low-dimensional PDEs can be solved effectively via traditional mesh-dependent methods such as finite difference methods [26], finite elements methods [21], finite volume methods [42], and spectral methods [3], developing efficient and accurate algorithmic frameworks for computing numerical solutions to high-dimensional PDEs remains an important and challenging topic due to the curse of dimensionality (i.e., the computational costs grow exponentially with respect to the dimensionality) [10].
Before solving a PDE, one must understand certain fundamental properties of the solution $u$ that follow from the PDE itself and from the data, namely, $f$ and $g$. These considerations align with the classical theory of PDEs [11], which focuses on establishing well-posedness (i.e., the existence and uniqueness of solutions) and their regularity and continuity. However, the underlying symbolic relationship among $u$, $f$ and $g$ often remains underexplored. To address this gap, we pose the following question:
Assuming that the PDE (1) admits an analytical solution $u$, and given the operators in $f$ and $g$, how are the operators in $u$ related to those in $f$ and $g$?
Understanding this symbolic connection is crucial for deriving fully interpretable, closed-form solutions, a feature lacking in most existing solution methods, particularly in the high-dimensional setting. As an illustrative example, we consider solving the PDE (1) via the least square method [9, 24, 37],which defines a straightforward loss to characterize the error of the estimated solution by
@@
$u^{*}=\mathrm{argmin}\left\{L(u)\;:\;u\in L^{2}(\Omega)\right\}.$ (3)
Because the search space $L^{2}(\Omega)$ is an infinite dimensional space, finding the solution $u^{*}$ is extremely challenging. Fortunately, the search space can be significantly simplified if one knows the operators appearing in the solution in advance, making the search of the optimal solution much easier and effective. This can be accomplished by adopting state-of-the-art symbolic machine learning approaches, including the efficient finite expression method (FEX) of [30].
To uncover the symbolic relationship, we leverage large language models (LLMs) as predictive tools. The process begins with the generation of a comprehensive, structured dataset composed of symbolic expressions derived from a diverse range of PDE types, including elliptic, parabolic, and hyperbolic equations, along with their associated boundary and initial conditions. These symbolic expressions are represented as computational trees (see Figure 5), where nodes are populated with unary (e.g., sin, cos, exp) and binary operators (e.g., +, -, *, /). This tree structure captures the hierarchical nature of mathematical expressions and ensures a systematic representation of the PDE solutions.
@@ 1.1 Related work
Artificial intelligence (AI) for Science takes advantage of modern advanced machine learning (ML) methods to facilitate, accelerate and enhance scientific discovery [40]. Its importance lies in the ability to handle massive datasets, uncover hidden patterns, and simulate complex phenomena. Inspired by the remarkable success of AI across various domains, deep neural network (DNN)-based approaches [25, 13] have become increasingly popular for solving (high-dimensional) PDEs. These methods offer certain appealing advantages over traditional numerical techniques, making them a compelling alternative for tackling complex problems. First, DNN-based approaches empirically mitigate the curse of dimensionality that plagues conventional discretization approaches, leading to more efficiency in handling high-dimensional PDEs [16]. Second, DNNs, acting as mesh-free universal function approximators [7, 19, 13], can be trained to learn highly complex PDE solutions without the need of prespecifying basis functions or meshes [1, 33]. Hence, they are highly suitable to approximate complex solution spaces that challenges the traditional mesh-based approaches. Third, the solutions represented by DNNs can be evaluated efficiently once trained, which turns out to be a crucial property for tasks, including uncertainty quantification and real-time optimal control, require calling PDE solutions repeatedly [46]. Last but not least, continuous progress in designing comprehensive neural network architectures, such as Fourier neural operators and advanced physics-informed neural network frameworks, alongside theoretical advancements in understanding their approximation capabilities, are making these methods increasingly applicable [31, 28].
@@
However, several critical factors often hinder deep neural networks (DNNs) from achieving highly accurate solutions, even for relatively simple problems. These include the need for large, diverse, and high-quality training data [41, 4], sensitivity to hyperparameter selection [2], the challenge of optimizing highly nonconvex objectives [8, 6], issues like vanishing and exploding gradients [12], and poor generalization capabilities [44]. Additionally, solutions produced by DNNs often lack interpretability, preventing users from leveraging insights for future decision-making [34]. To overcome these challenges and achieve both highly accurate and interpretable solutions for PDEs, [30] recently introduced the finite expression method (FEX). This approach leverages recent advances in approximation theory, which demonstrate that high-dimensional functions can be effectively approximated using functions composed of simple operators [35, 36, 45]. This methodology seeks approximate PDE solutions within the space of functions composed of finitely many analytic expressions, effectively avoiding the curse of dimensionality. In FEX, the mathematical expression representing the PDE solution is modeled as a binary tree, where each node holds either a unary or binary operator chosen from pre-specfied operator sets. The goal is to find the optimal sequence of operators and coupling parameters that minimize the objective function associated with the PDE. This leads to a combinatorial optimization (CO) problem, which is then solved by a deep reinforcement learning (RL) method with the policy gradient optimizer. The FEX outperforms certain existing numerical PDE solvers in terms of accuracy, interpretability and memory efficiency, making it a promising approach for solving high-dimensional PDEs and other complex problems; see e.g., [38, 39, 18, 29].
2 Theoretical Insight for Poisson Equations
While providing a complete answer to the question raised in the Introduction remains a significant challenge, in this section we take an initial step toward addressing it by offering theoretical insights that illuminate key aspects of the problem and motivate directions for future research.
Specifically, we focus on the classical Poisson equation with Dirichlet boundary condition
$-\Delta u=f\quad\text{in }\Omega,\qquad u=g\quad\text{on }\partial\Omega,$ (4)
where $\Omega\subseteq\mathbb{R}^{d}$ is a compact domain. By analyzing this simple class of elliptic PDEs under, we aim to shed light on the structural properties and solution behavior relevant to the broader question, thereby laying the groundwork for more comprehensive theoretical development in future work.
Theorem 1.
Let $\Omega\subset\mathbb{R}^{d}$ be a compact domain with boundary $\partial\Omega$ such that the distance function $\mathcal{D}(x):=\mathrm{dist}(x,\partial\Omega)$ admits an explicit analytical expression. Suppose the source term $f\in C^{0,\alpha}(\Omega)$ for some $\alpha\in(0,1]$ (i.e., $f$ is Hölder continuous), and the boundary condition $g$ is the restriction to $\partial\Omega$ of a function in $C^{2}(\overline{\Omega})$. Let $u$ be the analytical solution to the Possion equation (4). Then for any $\delta\in(0,1)$, there exists an approximating function $\tilde{u}:\mathbb{R}^{d}\to\mathbb{R}$ such that:
  1. 1.
    $\|u-\tilde{u}\|_{L^{2}(\Omega)}\leq\delta$,
  2. 2.
    $\tilde{u}$ is constructed using only the elementary arithmetic operators +,×$+,\;\times,\;(\cdot)^{2}$ and the operators appearing in the evaluation of $\mathcal{D}(\cdot)$, $f(\cdot)$, and $g(\cdot)$,
  3. 3.
    The total number of operations needed to evaluate $\tilde{u}(x)$ at any $x\in\mathbb{R}^{d}$ satisfies:
    $\texttt{ops}(\tilde{u})=\texttt{poly}(\delta^{-1})\cdot\left(\texttt{ops}(f)+\texttt{ops}(g)\right),$
    where $\texttt{poly}(\delta^{-1})$ denotes a quantity polynomial in $\delta^{-1}$.
Proof.
See Appendix B. ∎
This theorem offers a partial constructive resolution to the open research question posed in the previous section: it identifies the specific operators needed to represent a $\delta$-accurate approximation of the PDE solution $u$ in terms of the operators used to define the source term $f$, the boundary condition $g$, and the domain geometry via the distance function $\mathcal{D}$. The central insight is that an accurate approximation $\tilde{u}$ can be constructed using only a limited set of algebraic operations, along with the computational building blocks inherent in $f$, $g$, and $\mathcal{D}$. This result carries several important implications. (1) In contexts where one seeks symbolic models for physical phenomena, the theorem guarantees the existence of a low-complexity symbolic approximation of $u$ that relies solely on a structured and interpretable set of primitive operations. (2) It also informs the design of neural architectures for PDE approximation, suggesting that imposing operator priors, i.e., by reusing the operators present in $f$ and $g$, is sufficient to achieve arbitrarily high approximation accuracy, thereby promoting efficiency and interpretability. (3) The polynomial bound on computational complexity with respect to $\delta^{-1}$ underscores the practical feasibility and scalability of such approximations. Taken together, these insights lay a foundation for a broader theory of operator-preserving approximations for general PDEs, with significant potential impact in scientific machine learning, computational physics, and algorithmic model design.
@@ 3 Methodology
3.1 Binary Computational Trees for Expressions
Expressions in PDEs are represented as binary computational trees, as depicted in Figure 5 in Appendix A. These trees provide a structured and hierarchical way to encode mathematical expressions, enabling efficient exploration and manipulation of the solution space.
The construction of the tree proceeds in two main stages. Depth Specification: The maximum depth of the tree is specified in advance, determining the number of hierarchical levels and thereby controlling the expressive capacity of the resulting representations. While deeper trees enable the generation of more complex expressions, they also entail higher computational costs. Recursive Construction: Given the specified depth, the tree is constructed recursively. Starting from the root node, each node is expanded by generating child nodes according to predefined operator rules.
Each node in the computational tree encodes an operator, which may be either a unary or binary operator. Unary Operators: Unary operators act on a single operand and include a variety of nonlinear transformations such as trigonometric functions ($\sin(\cdot)$, $\cos(\cdot)$, $\tan(\cdot)$), exponential functions ($\exp(\cdot)$), logarithmic functions ($\lg(\cdot)$, $\ln(\cdot)$), and other elementary operations (e.g., $\sqrt{\cdot}$, ||$\cdot$|| ⋅ |, and power functions $(\cdot)^{k}$ for $k\in\mathbb{Z}$). Each unary operator node is augmented with two scalar parameters, $\alpha$ and $\beta$, which enable the composition of the operator with an affine transformation, yielding expressions of the form $\alpha\cdot u(\cdot)+\beta$, where $u(\cdot)$ denotes the unary operation. This parameterization allows the tree to flexibly represent both nonlinear and linearly adjusted transformations, thereby enhancing its expressiveness. Binary Operators: Binary operators operate on two operands and include the standard arithmetic operations: addition ($+$), subtraction (--), multiplication ($\times$), and division (///). Each binary operator node combines the outputs of its left and right child subtrees to produce more complex expressions. For example, a multiplication node may combine two distinct sub-expressions, thereby enabling the construction of richer symbolic forms through hierarchical composition.
@@ 3.2 Data Generation for Equation Types
We next describe the data generation process for constructing a structured dataset of symbolic expressions derived from various PDE types. The data generation process is outlined in Figure 1.
@@
Specifically, the construction of a single data point consists of the following steps: PDE Type Specification. We begin by specifying the type of PDE and the type of boundary condition, based on the associated differential and boundary operators $D$ and $B$. This information forms the first component of the data point, identifying the PDE and boundary condition types. Generation of Random Function $u$. Next, we construct a random solution function $u$ using a binary computational tree of predefined depth, as described in Section 3.1. The tree is populated with randomly selected unary and binary operators from a sufficiently expressive operator set. This yields a symbolic expression for $u$, from which we compute the corresponding right-hand-side function $f$ and boundary function $g$, together forming the second component of the data point. Postfix Representation. To facilitate processing by large language models (LLMs), the expressions for $u$, $f$, and $g$ are converted into postfix notation. This format simplifies tokenization, avoids the use of parentheses, and explicitly encodes operator precedence through position. Importantly, it captures the structural relationships among operators, which helps LLMs learn symbolic patterns. The postfix expressions of $f$ and $g$ constitute the model input, while the postfix form of $u(x)$ serves as the output label.
3.3 Fine-tuning LLM
With a generated data set, we can then fine-tune a certain LLM to learn the symbolic relastion in PDEs. The whole pipeline is summarized in Figure 2, and we will describe each component in detail for the rest of this subsection.
Figure 2: Overview of the fine-tuning pipeline
Model Selection. Encoder-decoder and decoder-only architectures are two fundamental types of LLMs, each with distinct design principles and strengths. Encoder-decoder models (e.g., T5 [32] and BART [27]) consist of an encoder that transforms the entire input sequence into a rich latent representation, and a decoder that generates the output sequence token by token, attending to the encoder’s outputs via cross-attention. This architecture is well-suited for tasks such as machine translation, summarization, and question answering, where understanding the full input before generating the output is crucial. In contrast, decoder-only models (e.g., Llama3 [14]) treat the input and output as a single continuous sequence and generate tokens autoregressively using causal self-attention, meaning each token can only attend to previous tokens. This design simplifies the model architecture and training process, and has proven particularly effective for open-ended context generation and few-shot prompting. Based on our empirical results, decoder-only models are preferable for our task, as they more accurately predict the symbolic operators associated with PDE solutions.
Tokenization. We next perform tokenization to convert raw text into tokens that the model can process numerically, enabling efficient representation and learning. Specifically, each data point (in raw text) is serialized into a prompt followed by its target operator sequence, as demonstrated in the following.
$\underbrace{\texttt{Type:}<\mathrm{PDE}>\,\texttt{|}\,\texttt{RHS:}<\mathrm{ops}>\,\texttt{|}\,\texttt{BC\_Type:}<\mathrm{bc}>\,\texttt{|}\,\texttt{Solution:}}_{\text{PROMPT}}\;\|\|\underbrace{\texttt{<op}_{1}\ \dots\ \texttt{op}_{n}\,\texttt{<EOS>}\texttt{>}}_{\text{TARGET}}$
Next, a pretrained tokenizer is applied to convert the input text into its corresponding token sequence. In this work, we use the SentencePiece  [^1] Available at https://github.com/google/sentencepiece. tokenizer and explicitly set the PAD token to match the EOS token, ensuring that padding positions are never treated as valid outputs. To accommodate memory constraints, we truncate the prompt to a maximum of 256 tokens and the target to a maximum of 64 tokens, resulting in a concatenated sequence length of 300. Additionally, we construct a binary attention mask, where real tokens are marked with 1 and padding positions with 0. This mask is supplied to the model to ensure that self-attention computations ignore padded positions.
Fine-tuning. Given a pretrained decoder-only language model and a tokenized dataset, we proceed to the fine-tuning phase, where the model’s parameters are updated via stochastic gradient-based optimization. Specifically, we minimize the cross-entropy loss between the model’s predicted token distributions and the ground truth tokens in the training data. This objective encourages the model to assign higher probabilities to correct next-token predictions, thereby adapting its knowledge to the target task or domain. For optimization, we use the AdamW optimizer, a variant of Adam [22] that decouples weight decay from the gradient updates, which has been shown to improve generalization performance in transformer-based models. To further enhance the efficiency of training, we adopt several computational techniques. First, we employ mixed-precision training, which allows computations to be performed in lower precision (e.g., float16) while maintaining a decent level of stability and accuracy. This significantly reduces memory usage and accelerates training. Second, we leverage Low-Rank Adaptation (LoRA) [20], a parameter-efficient fine-tuning technique that injects trainable low-rank matrices into specific layers (typically attention or feedforward layers) of the pretrained model while keeping the original model weights frozen. This approach greatly reduces the number of trainable parameters and memory footprint during training, making it especially suitable for large-scale models and resource-constrained environments. Together, these strategies enable effective and efficient fine-tuning of large language models on our tasks.